The smaller of two numbers is 10 less than 4 times the greater number. The greater number is 19 more than the smaller number. Find the 2 numbers.
Answer by Flannery(124)

Let x = the greater number.
4x-10=the smaller number.
x=the smaller number+19.
x=4x-10+19
x=4x+9
-3x=9
x=-3
check
4(-3)-10=the smaller number
-12-10=the smaller number
-22=the smaller number
if -3 is the greater number and -22 is the smaller number then -3=-22+19
-3=-22+19
-3=-3
so -3 is the greater number and -22 is the smaller number
Q.E.D.